# Stepwise Guided Policy Optimization: Coloring your Incorrect Reasoning in GRPO

### v4

@@ Abstract
Reinforcement learning (RL) has proven effective in strengthening the reasoning capabilities of large language models (LLMs). A widely adopted method, Group Relative Policy Optimization (GRPO) (Shao et al., 2024), has shown strong empirical results in training DeepSeek-R1 (Guo et al., 2025a). However, GRPO fails to update the policy when all responses within a group are incorrect (i.e., all-negative-sample groups). This limitation underscores a key gap between artificial and human intelligence: unlike humans, who can learn from mistakes, GRPO discards these signals. Our first contribution is to introduce a simple framework that mitigates the all-negative-sample issue by incorporating response diversity within groups using a step-wise judge model, which can be either directly trained or adapted from existing LLMs. We prove that this diversification can accelerate GRPO’s learning dynamics in a simplified setting. We also empirically validate the proposed stepwise guided policy optimization (SGPO) method, demonstrating consistent gains across model sizes (7B, 14B, 32B) in offline and online training on 9 benchmarks, including base and distilled variants. Our results highlight two advantages: (i) SGPO surpasses GRPO, especially in the early and mid-training stages where all-negative-sample groups are prevalent; and (ii) SGPO does not require judge models to generate correct answers, differentiating it from knowledge distillation methods.
@@ 1 Introduction
The rise of OpenAI-o1 (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025a), and Kimi-1.5 (Team et al., 2025) has highlighted the emergence of large AI reasoning models. Unlike instruction-tuned models (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023; Achiam et al., 2023), which produce quick responses by statistically inferring the next token, these new reasoning models deliberately decompose complex prompts (e.g., mathematical problems) into intermediate steps and work through chain-of-thought reasoning (Wei et al., 2022; Yao et al., 2023; Besta et al., 2024; Xiang et al., 2025). This slower yet more rigorous process yields greater accuracy and makes them more human-like, enabling success on more complex and challenging tasks (Yang et al., 2018; Shi et al., 2024; Jain et al., 2025). As generative AI applications move beyond simple conversational interfaces, the reasoning models are poised to become increasingly powerful and widely adopted in practice.
At the heart of this revolution lies post-training with outcome-based and verifiable rewards (Cobbe et al., 2021; Uesato et al., 2022; Zelikman et al., 2022; Singh et al., 2023; Hosseini et al., 2024; Lightman et al., 2024; Wang et al., 2024; Setlur et al., 2025; Zhang et al., 2025b), together with reinforcement learning (RL) methods (Schulman et al., 2015, 2017; Li et al., 2024b; Ahmadian et al., 2024; Shao et al., 2024; Xiong et al., 2025a), appreciated for their simplicity, intuitiveness, and practicality. A leading approach is proximal policy optimization (PPO) (Schulman et al., 2017), which relies on a critic (or value) model to estimate advantages. While essential in general RL tasks, this critic is often unnecessary in large language models (LLMs) due to their deterministic transition dynamics (Li et al., 2024b). This observation has inspired alternatives such as group relative policy optimization (GRPO) (Shao et al., 2024) and its extensions (Yu et al., 2025b; Liu et al., 2025b; Chu et al., 2025; Zhang et al., 2025a), which estimate advantages directly in a group-relative fashion.
@@
The key insight is that many reasoning tasks possess a structure where step-level correctness can be explicitly defined. This motivates the use of a step-wise judge model that evaluates trajectories by labeling each step as correct (1) or incorrect (0). Such a model can be trained directly (Xiong et al., 2025b) or adapted from existing LLMs (Zha et al., 2025; He et al., 2025) [^2] We do not have access to their judge models as it’s not publicly released, so we adapt our own from existing LLMs. . By grounding rewards in step-level correctness rather than speculative value estimates, our method mitigates reward hacking and yields clearer signals. Intuitively, this allows negative samples to be differentiated through their trajectories: while early-stage reasoning trajectories are of low-quality, these remain informative – much like partial credit in education, where intermediate steps still guide learning.
Our approach enables a holistic evaluation of multi-step reasoning by transforming negative samples from binary outcome rewards into graded, step-level rewards. Consider a negative sample with five reasoning steps $(a_{1},a_{2},a_{3},a_{4},a_{5})$. If the first error occurs at $a_{3}$, then $a_{1}$ and $a_{2}$ are correct, yielding a correctness proportion of $\tfrac{2}{5}$. To improve reliability, we adopt a Grok4-Heavy -inspired strategy where multiple independent judgments are obtained from the judge model, and the error position is determined by the majority vote. We further introduce two scaling parameters $\beta$ and $\gamma$ to downweight noisy or unreliable signals (see Eq. 2). Unlike PRMs, our approach avoids memory overhead and does not require costly step-level human annotations, thereby accelerating training.
@@
    We conduct the experiments demonstrating the effectiveness of our approach in improving LLM reasoning. Evaluations are undertaken across various model sizes (7B, 14B, 32B) in both offline and online settings with nine benchmarks, including base and distilled variants. Our results reveal two key benefits: (i) SGPO delivers improvements beyond the reach of GRPO, especially in the early and mid-stages of training where all-negative-sample groups are common; (ii) SGPO does not rely on more powerful judge models generating correct answers, allowing it to be distinguish from knowledge distillation methods.
@@ 2 Preliminaries and Technical Background
Modern LLMs are built based on the Transformer architecture (Vaswani et al., 2017) and generate responses $\mathbf{y}=(a_{1},\ldots,a_{H})$ to user prompts $\mathbf{x}$, where each token $a_{h}\in\mathcal{V}^{\star}$, with $\mathcal{V}$ denoting the vocabulary and $\mathcal{V}^{\star}$ the set of all possible token sequences. We view the LLM as a policy $\pi_{\theta}(\mathbf{y}|\mathbf{x})$ parameterized by $\theta$, assigning probabilities to responses $\mathbf{y}$ given $\mathbf{x}$. The policy operates in an auto-regressive way:
@@
$J(\theta)={\mathbb{E}}_{\mathbf{x}\sim\rho,\mathbf{y}\sim\pi_{\theta_{\textnormal{old}}}(\cdot|\mathbf{x})}\left[\min\left\{\tfrac{\pi_{\theta}(\mathbf{y}|\mathbf{x})}{\pi_{\theta_{\textnormal{old}}}(\mathbf{y}|\mathbf{x})}r(\mathbf{x},\mathbf{y}),\texttt{clip}\left\{\tfrac{\pi_{\theta}(\mathbf{y}|\mathbf{x})}{\pi_{\theta_{\textnormal{old}}}(\mathbf{y}|\mathbf{x})},1-\epsilon,1+\epsilon\right\}r(\mathbf{x},\mathbf{y})\right\}\right].$
The group relative policy optimization (GRPO) and its variants (Yu et al., 2025b; Liu et al., 2025b; Chu et al., 2025; Zhang et al., 2025a) adopt this framework but estimate gradients using groups of samples. For each prompt $\mathbf{x}$, GRPO samples responses $\mathbf{y}_{1},\ldots,\mathbf{y}_{G}$} from $\pi_{\theta_{\textnormal{old}}}$ and maximizes the objective function in the form of
$J(\theta)=\tfrac{1}{G}\sum_{i=1}^{G}\left[\min\left\{\tfrac{\pi_{\theta}(\mathbf{y}_{i}|\mathbf{x})}{\pi_{\theta_{\textnormal{old}}}(\mathbf{y}_{i}|\mathbf{x})}A_{i},\texttt{clip}\left\{\tfrac{\pi_{\theta}(\mathbf{y}_{i}|\mathbf{x})}{\pi_{\theta_{\textnormal{old}}}(\mathbf{y}_{i}|\mathbf{x})},1-\epsilon,1+\epsilon\right\}A_{i}\right\}\right],$
@@ 3.1 A Step-wise Judge Model
In our experiment, we adapt the judge model from existing LLMs, either closed-source (o4-mini, Claude3.7) or open-source (DeepSeek-V3-0324, Qwen3-235B-A22B, QwQ-32B). To enhance reliability and further reduce variance in the reward signal, we employ the following protocol: (i) alongside the candidate response, we provide a reference solution drawn from a supervised fine-tuning dataset with correct answers and reasoning trajectories, anchoring the intended solution path and enabling error localization; and (ii) we elicit step-wise evaluation rather than holistic evaluation. The judge model justifies correctness or flags an error sentence by sentence, identifies the first clear mistake, and then traces how this error propagates to the final incorrect conclusion.
@@
$r_{\texttt{SGPO}}(\mathbf{y})=\left\{\begin{array}[]{cl}1,&\textnormal{if the final answer of }\mathbf{y}\textnormal{ is correct},\\ \frac{1}{1+\exp(\beta(\texttt{RTS}(\mathbf{y})-\gamma))},&\textnormal{otherwise}.\end{array}\right$. (2)
where $\gamma>0$ and $\beta>0$ are two parameters to decide scale threshold and scale intensity, respectively. This design ensures that the model receives a more informative gradient signal during training, thereby encouraging refinement of partially correct reasoning rather than indiscriminate penalization of all incorrect outputs. This specification of $r_{\texttt{SGPO}}$ can be directly incorporated into the advantage calculation in Eq. 1. As a consequence, we refer to SGPO as GRPO using Eq. 2.
@@
Our approach differs from knowledge distillation (e.g. Kang et al., 2023; Gu et al., 2024). The student model trained via knowledge distillation inherits the judge model’s failure, since it only imitates the judge model’s outputs. For instance, consider the AIME problem: “The twelve letters $\{A,B,C,D,E,F,G,H,I,J,K,L\}$ are randomly grouped into six pairs. Each pair is ordered alphabetically to form a two-letter word, and the six words are listed alphabetically, such as $AB,CD,EF,GH,IJ,KL$. The probability that the last word listed contains $G$ is $\frac{m}{n}$ with $m,n$ coprime. Find m+nm+n”. Neither the student model (DeepSeek-R1-Distill-Qwen-7B) nor the judge models (DeepSeek-V3-0324) can solve this problem within 16 rollouts. In contrast, SGPO leverages the judge model to identify mistakes in the student’s reasoning, providing learning signals that go beyond imitation and enabling improvements unattainable by knowledge distillation.
@@ 3.2 Accelerating Learning Dynamics
We present a theoretical analysis to explain why SGPO outperforms GRPO. To this end, we consider a simplified setting with a reasoning horizon of $H=2$, where each step admits two possible actions $a_{h}\in{1,2}$ for $h=1,2$. This configuration follows prior works (Dayan, 1991; Li et al., 2024b), where analogous examples were employed to validate theoretical insights. Without loss of generality, we assume a unique ground-truth response $\mathbf{y}_{\mathbf{x}}^{\star}=(2,2)$ for the prompt $\mathbf{x}$. The algorithm iteratively updates the policy parameter $\theta$ using samples drawn from the current policy $\pi_{\theta}$. For clarity, we restrict the sample space to (1,1),(2,1),(2,2){(1,1),(2,1),(2,2)}, excluding $(1,2)$ since a correct reasoning step is unlikely to, and should not, follow an incorrect precursor.
@@
To illustrate the effect of SPO, we analyze the learning dynamics of SGPO and GRPO in this simplified setting. Under GRPO, the rewards are assigned as $r((2,2))=1$ and $r((2,1))=r((1,1))=0$, meaning that only selecting the “good” action 2 at both steps yields a positive reward. In contrast, SGPO assigns $r_{\texttt{SGPO}}((2,2))=1$, $r_{\texttt{SGPO}}((2,1))=\tfrac{1}{2}$ and $r_{\texttt{SGPO}}((1,1))=0$. The difference is that partial progress – choosing the “good” action 2 in the first step but failing at the second – receives no credit in GRPO yet proportional credit in SGPO. Here, $\tfrac{1}{2}$ is chosen for illustrative purposes to convey the qualitative behavior of the reward mechanism, while the exact values used in experiments are determined by Eq. 2.
Figure 1: SGPO outperforms GRPO in learning a “good” action in the first step.
In our analysis, we examine the population-level learning dynamics with $G=2$, omitting clipping and importance sampling. Let $p_{\texttt{GRPO}}^{(k)}$ and $q_{\texttt{GRPO}}^{(k)}$ denote the probability of selecting the “good” action in the first step at iteration $k$ under GRPO, and that of selecting the “good” action in the second step conditioned on a correct first step. Analogously, $p_{\texttt{SGPO}}^{(k)}$ and $q_{\texttt{SGPO}}^{(k)}$ denote the corresponding probabilities under SGPO. Our theoretical findings are summarized in the following theorem.

### v5

@@ Abstract
Reinforcement learning (RL) has proven effective in strengthening the reasoning capabilities of large language models (LLMs). A widely adopted method, Group Relative Policy Optimization (GRPO) (Shao et al., 2024), has shown strong empirical results in training recent reasoning models (Guo et al., 2025a), but it fails to update the policy when all responses within a group are incorrect (i.e., all-negative-sample groups). This limitation highlights a gap between artificial and human intelligence: unlike humans, who can learn from mistakes, GRPO discards these failure signals. We introduce a simple framework to mitigate the all-negative-sample issue by incorporating response diversity within groups using a step-wise judge model, which can be trained directly or adapted from existing LLMs. In a simplified setting, we prove that this diversification accelerates GRPO’s learning dynamics. We then empirically validate Stepwise Guided Policy Optimization (SGPO) across model sizes (7B, 14B, 32B) in both offline and online training on nine reasoning benchmarks (including base and distilled variants). Overall, SGPO improves average performance and is effective in early and mid-training when all-negative groups are prevalent, while improvements are not uniform across every benchmark and depend on the structure and informativeness of negative samples. Finally, SGPO does not require the judge model to generate correct solutions, distinguishing it from knowledge distillation methods.
Figure 1: Main pipeline for Stepwise Guided Policy Optimization. On each llama, $r$ indicates the reward of the sampled response and AA indicates response’s advantage through group relative computation.
@@ 1 Introduction
The rise of OpenAI-o1 (Jaech et al., 2024), DeepSeek-R1 (Guo et al., 2025a), and Kimi-1.5 (Team et al., 2025) has highlighted the emergence of large AI reasoning models. Unlike instruction-tuned models (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023; Achiam et al., 2023), which produce quick responses by statistically inferring the next token, these new reasoning models deliberately decompose complex prompts (e.g., mathematical problems) into intermediate steps and work through chain-of-thought reasoning (Wei et al., 2022; Yao et al., 2023; Besta et al., 2024; Xiang et al., 2025). This slower yet more rigorous process yields greater accuracy and makes them more human-like, enabling success on more complex and challenging tasks (Yang et al., 2018; Shi et al., 2024; Jain et al., 2025). As generative AI applications move beyond single-turn chat and question-answering, these reasoning models are poised to become more powerful and widely adopted, positioning them as a foundational component of modern AI systems.
At the heart of this revolution lies post-training with outcome-based and verifiable rewards (Cobbe et al., 2021; Uesato et al., 2022; Zelikman et al., 2022; Singh et al., 2023; Hosseini et al., 2024; Lightman et al., 2024; Wang et al., 2024; Setlur et al., 2025; Zhang et al., 2025b), together with reinforcement learning (RL) methods (Schulman et al., 2015, 2017; Li et al., 2024b; Ahmadian et al., 2024; Shao et al., 2024; Xiong et al., 2025a), appreciated for their simplicity, intuitiveness, and practicality. A leading approach is proximal policy optimization (PPO) (Schulman et al., 2017), which relies on a critic (or value) model to estimate advantages. While essential in general RL tasks, this critic is often unnecessary in large language models (LLMs) due to their deterministic transition dynamics (Li et al., 2024b). This observation has inspired alternatives such as group relative policy optimization (GRPO) (Shao et al., 2024) and its extensions (Yu et al., 2025b; Liu et al., 2025b; Chu et al., 2025; Zhang et al., 2025a), which estimate advantages directly in a group-relative fashion (normalizing rewards across multiple samples for the same prompt).
@@
A common observation is that many reasoning tasks possess a structure where step-level correctness can be explicitly defined. This motivates the use of a step-wise judge model that evaluates trajectories by labeling each step as correct (1) or incorrect (0). Such a model can be trained directly (Xiong et al., 2025b) or adapted from existing LLMs (Zha et al., 2025; He et al., 2025) [^2] We do not have access to their judge models as it’s not publicly released, so we adapt our own from existing LLMs. . By grounding rewards in step-level correctness rather than speculative value estimates, our method mitigates reward hacking and yields clearer signals. Intuitively, this allows negative samples to be differentiated through their trajectories: while early-stage reasoning trajectories are of low-quality, these remain informative – much like partial credit in education, where intermediate steps still guide learning.
Our approach enables a holistic evaluation of multi-step reasoning by transforming negative samples from binary outcome rewards into graded, step-level rewards. Consider a negative sample with five reasoning steps $(a_{1},a_{2},a_{3},a_{4},a_{5})$. If the first error occurs at $a_{3}$, then $a_{1}$ and $a_{2}$ are correct, yielding a correctness proportion of $\tfrac{2}{5}$. To improve reliability, we adopt a Grok4-Heavy -inspired strategy where multiple independent judgments are obtained from the judge model, and the error position is determined by the majority vote. We further introduce two scaling parameters $\beta$ and $\gamma$ to downweight noisy or unreliable signals (see Eq. 2). Unlike PRMs, our approach avoids memory overhead and does not require costly step-level human annotations, thereby accelerating training. In this work, we focus on outcome-based post-training with group-relative updates (GRPO-style) for structured reasoning tasks; extending our approach to arbitrary reward settings is beyond the scope of this paper and left to future work.
@@
    We conduct experiments demonstrating the effectiveness of our approach in improving LLM reasoning. Evaluations are undertaken across various model sizes (7B, 14B, 32B) in both offline and online settings with nine benchmarks, including base and distilled variants, using GRPO as the primary baseline given our focus on outcome-based group-relative RLVR. Our results reveal two key benefits: (i) SGPO delivers improvements beyond the reach of GRPO, especially in the early and mid-stages of training where all-negative-sample groups are common; (ii) SGPO does not rely on more powerful judge models generating correct answers, allowing it to be distinguished from knowledge distillation.
@@ 2 Preliminaries and Technical Background
LLM finetuning.
LLM finetuning typically consists of pre-training and post-training. Pre-training equips the model with broad language understanding and generation capabilities, while post-training adapts the model to specific downstream objectives (e.g., improving mathematical problem solving via reasoning). In post-training, a model usually first undergoes imitation learning (e.g., supervised finetuning on human or expert trajectories, or direct distillation from stronger models), and then further improves by training on self-generated responses paired with feedback indicating whether the responses are good or bad. Two common feedback-driven paradigms are reinforcement learning from human feedback (RLHF), where a learned reward model (trained from human preferences) assigns rewards to model outputs, and reinforcement learning from verifiable rewards (RLVR), where rewards are computed by an automatic verifier (often used for math problems with checkable answers). We introduce GRPO, a representative RLVR method, below.
Policy optimization.
Modern LLMs are built based on the Transformer architecture (Vaswani et al., 2017) and generate responses $\mathbf{y}=(a_{1},\ldots,a_{H})$ to user prompts $\mathbf{x}$, where each token $a_{h}\in\mathcal{V}^{\star}$, with $\mathcal{V}$ denoting the vocabulary and $\mathcal{V}^{\star}$ the set of all possible token sequences. We view the LLM as a policy $\pi_{\theta}(\mathbf{y}|\mathbf{x})$ parameterized by $\theta$, assigning probabilities to responses $\mathbf{y}$ given $\mathbf{x}$. The policy operates in an auto-regressive way as follows (Agarwal et al., 2020; Mei et al., 2021; Li et al., 2024b):
@@
$J(\theta)={\mathbb{E}}_{\mathbf{x}\sim\rho,\mathbf{y}\sim\pi_{\theta_{\textnormal{old}}}(\cdot|\mathbf{x})}\left[\min\left\{\tfrac{\pi_{\theta}(\mathbf{y}|\mathbf{x})}{\pi_{\theta_{\textnormal{old}}}(\mathbf{y}|\mathbf{x})}r(\mathbf{x},\mathbf{y}),\texttt{clip}\left(\tfrac{\pi_{\theta}(\mathbf{y}|\mathbf{x})}{\pi_{\theta_{\textnormal{old}}}(\mathbf{y}|\mathbf{x})},1-\epsilon,1+\epsilon\right)r(\mathbf{x},\mathbf{y})\right\}\right],$
where $\texttt{clip}(x,1-\varepsilon,1+\varepsilon):=\max\{\min\{x,1+\varepsilon\},1-\varepsilon\}$. The group relative policy optimization (GRPO) and its variants (Yu et al., 2025b; Liu et al., 2025b; Chu et al., 2025; Zhang et al., 2025a) adopt this framework but estimate gradients using groups of samples. For each prompt $\mathbf{x}$, GRPO samples responses $\mathbf{y}_{1},\ldots,\mathbf{y}_{G}$} from $\pi_{\theta_{\textnormal{old}}}$. We aim at maximizing the objective function in the form of
$J(\theta)={\mathbb{E}}_{\mathbf{x}\sim\rho,\{\mathbf{y}_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\textnormal{old}}}(\cdot|\mathbf{x})}\left[\tfrac{1}{G}\sum_{i=1}^{G}\min\left\{\tfrac{\pi_{\theta}(\mathbf{y}_{i}|\mathbf{x})}{\pi_{\theta_{\textnormal{old}}}(\mathbf{y}_{i}|\mathbf{x})}A_{i},\texttt{clip}\left(\tfrac{\pi_{\theta}(\mathbf{y}_{i}|\mathbf{x})}{\pi_{\theta_{\textnormal{old}}}(\mathbf{y}_{i}|\mathbf{x})},1-\epsilon,1+\epsilon\right)A_{i}\right\}\right],$
@@ 3.1 A Step-wise Judge Model
In our experiment, we adapt the judge model from existing LLMs, either closed-source (o4-mini, Claude3.7) or open-source (DeepSeek-V3-0324, Qwen3-235B-A22B, QwQ-32B). To enhance reliability and further reduce variance in the reward signal, we employ the following protocol: (i) alongside the candidate response, we provide the judge with a reference solution (e.g., a gold final answer, a brief solution outline, or a full reasoning trace when available); in our experiments, we draw this solution from a supervised fine-tuning dataset with correct answers and reasoning trajectories, which anchors the intended solution path and enables error localization; and (ii) we elicit step-wise evaluation rather than holistic evaluation. The judge model justifies correctness or flags an error sentence by sentence, identifies the first clear mistake, and then traces how this error propagates to the final incorrect conclusion.
@@
$r_{\texttt{SGPO}}(\mathbf{y})=\left\{\begin{array}[]{cl}1,&\textnormal{if the final answer of }\mathbf{y}\textnormal{ is correct},\\ \frac{1}{1+\exp(-\beta(\texttt{RTS}(\mathbf{y})-\gamma))},&\textnormal{otherwise}.\end{array}\right$. (2)
where $\gamma>0$ and $\beta>0$ (taken as 10 and 0.5 in the actual implementation) are two parameters to decide scale threshold and scale intensity, respectively. This design ensures that the model receives a more informative gradient signal during training, thereby encouraging refinement of partially correct reasoning rather than indiscriminate penalization of all incorrect outputs. This specification of $r_{\texttt{SGPO}}$ can be directly incorporated into the advantage calculation in Eq. 1. As a consequence, SGPO keeps the same rollout pipeline as GRPO and the same outcome-based supervision, and only replaces the reward used in within-group advantage computation from $r(x,y)$ to $r_{\texttt{SGPO}}(y)$ in Eq. 2.
@@
Our approach differs from knowledge distillation (e.g. Kang et al., 2023; Gu et al., 2024). The student model trained via knowledge distillation inherits the judge the model’s failure, since it only imitates the judge model’s outputs. In contrast, SGPO leverages the judge model to identify mistakes in the student’s reasoning, providing learning signals that go beyond imitation and enabling improvements unattainable by knowledge distillation.
@@ 3.2 Accelerating Learning Dynamics
We present a theoretical analysis to explain why SGPO outperforms GRPO. We study an idealized setting in which the step-wise judge provides accurate first-error localization. Even in this regime, establishing a general separation is technically subtle, so we focus on a stylized example. To this end, we consider the case when $H=2$, where each step admits two possible actions $a_{h}\in{1,2}$ for $h=1,2$. Extending the analysis to general horizons or action spaces and imperfect (noisy or biased) judges is left to future work. This configuration follows prior works (Dayan, 1991; Li et al., 2024b), in which analogous examples were employed to validate theoretical insights. Without loss of generality, we assume a unique ground-truth response $\mathbf{y}_{\mathbf{x}}^{\star}=(2,2)$ for the prompt $\mathbf{x}$. For clarity, we restrict the sample space to (1,1),(2,1),(2,2){(1,1),(2,1),(2,2)}, excluding $(1,2)$ since a correct reasoning step is unlikely to, and should not, follow an incorrect precursor.
@@
To illustrate the effect of SGPO, we analyze the learning dynamics of SGPO and GRPO in this simplified setting. Under GRPO, the rewards are assigned as $r((2,2))=1$ and $r((2,1))=r((1,1))=0$, meaning that only selecting the “good” action 2 at both steps yields a positive reward. In contrast, SGPO assigns $r_{\texttt{SGPO}}((2,2))=1$, $r_{\texttt{SGPO}}((2,1))=\tfrac{1}{2}$ and $r_{\texttt{SGPO}}((1,1))=0$. The difference is that partial progress – choosing the “good” action 2 in the first step but failing at the second – receives no credit in GRPO yet proportional credit in SGPO. Here, $\tfrac{1}{2}$ is chosen for illustrative purposes to convey the qualitative behavior of the reward mechanism, while the exact values used in experiments are determined by Eq. 2.
The algorithm iteratively updates the policy parameter $\theta$ using samples drawn from the current policy $\pi_{\theta}$. We rewrite the generic GRPO update with a step size $\eta>0$ as follows,
where $N$ is the number of prompts, $G$ is the number of groups, $H$ is the number of reasoning steps in each response, $s_{\theta}(\mathbf{x}^{i},a^{i,k}_{1:h-1}):=\nabla\theta\log\pi_{\theta}(a_{t}|\mathbf{x},a_{1:h-1})$ is the score function, and the advantage $A_{i,k}$ is defined by Eq. 1 for each question $\mathbf{x}^{i}$. To distinguish, we denote $g_{\texttt{GRPO}}(\cdot)$ as the gradient estimator using classical outcome reward model $r$, and $g_{\texttt{SGPO}}(\cdot)$ as the gradient estimator using the reward $r_{\texttt{SGPO}}$ as proposed in Section 3.1.
In our analysis, we examine the population-level learning dynamics with $G=2$, omitting clipping and importance sampling. In practice, importance sampling and clipping are important for training stability; we omit them here to simplify the analysis and leave their theoretical treatment to future work. For simplicity, we perform our analysis in the likelihood space rather than in the parameter space directly. Indeed, we define the key quantities as follows
$p\doteq\pi_{\theta_{1}}(a_{1}=2,|,\mathbf{x})=\tfrac{e^{\theta_{1}^{\mathbf{x},2}}}{e^{\theta_{1}^{\mathbf{x},1}}+e^{\theta_{1}^{\mathbf{x},2}}},\qquad q\doteq\pi_{\theta_{2}}(a_{2}=2,|,\mathbf{x},a_{1}=2)=\tfrac{e^{\theta_{2}^{\mathbf{x},2,2}}}{e^{\theta_{2}^{\mathbf{x},2,1}}+e^{\theta_{2}^{\mathbf{x},2,2}}}.$
and
The responses can be drawn i.i.d. from the distribution as follows,
The SGPO and GRPO training dynamics with population-level policy gradient can be computed exactly for the stylized model as follows,
Since $g_{\texttt{GRPO}}(\theta)$ and $g_{\texttt{SGPO}}(\theta)$ concentrate around $\bar{g}_{\texttt{GRPO}}(\theta)$ and $\bar{g}_{\texttt{SGPO}}(\theta)$ when the number of samples in each group is sufficiently large, it is reasonable to analyze the population-level dynamics for illustration. Note that the high-probability guarantees for the sample-level dynamics can be derived using concentration inequalities under certain conditions.
Denote $p_{\texttt{GRPO}}^{(k)}$ and $q_{\texttt{GRPO}}^{(k)}$ as the value of quantity $p$ and $q$ at iteration $k$ under GRPO. Analogously, $p_{\texttt{SGPO}}^{(k)}$ and $q_{\texttt{SGPO}}^{(k)}$ are the corresponding probabilities $p$ and $q$ at iteration $k$ under SGPO. We can explicitly write down the SGPO and GRPO update rules with $\eta=1$ as follows,
where the functions $f_{ij}$ are defined by
$f_{11}(p)=\log(p)+p(1-p)-\log(1-p+pe^{p(1-p)}),$ (4)
$f_{21}(p,q)=\log(p)+p(1-p)q-\log(1-p+pe^{p(1-p)q}),$
$f_{12}(p,q)=\log(q)+p^{2}q(1-q)-\log(1-q+qe^{p^{2}q(1-q)}),$
$f_{22}(p,q)=\log(q)+pq(1-q)-\log(1-q+qe^{pq(1-q)}).$
Our theoretical findings are summarized in the following theorem.